Determine if the system has 1 solution, no solutions, or zero solutions:
5x-15y=35
15y-10x=-20

Answer by jim_thompson5910(35256)

5x-15y=35
15y-10x=-20

is the same as

5x-15y=35
-10x+15y=-20

Add the two equations to get


5x-15y=35
-10x+15y=-20
-------------
-5x+0y = 15

So -5x = 15

This means that we can solve for x to get a specific number. We then use that solution for x to solve for y.

Overall, this system has 1 solution.